OpenStudy (anonymous):

find the exact soltions, in terms of radians, 2 cos (3theta) + square root of 3 = 0

OpenStudy (anonymous):

\[ 2\cos{3\theta} = -\sqrt3 \implies \cos{3\theta} = -\frac{\sqrt3}{2} \] Can you solve it beyond this point?